OpenStudy (anonymous):

Find the 110th term of the sequence –7 , 3, 13, 23,... Help me please

OpenStudy (anonymous):

i dont think thats right...

OpenStudy (anonymous):

i just need to know how to find the answer. like the formula for something

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

Sorry, I was helping someone else with a similar problem and I was using their data. Sorry again. -7,3,13,23 This is an arithmetic sequence since there is a common difference between each term. In this case, adding 10 to the previous term in the sequence gives the next term. In other words, an=an-1+d. Arithmetic Sequence: d=10 Reorder the polynomial n-1 alphabetically from left to right, starting with the highest order term. an=-7+10(n-1) Multiply 10 by each term inside the parentheses. an=-7+(10n-10) Subtract 10 from -7 to get -17. an=-17+10n Reorder the polynomial -17+10n alphabetically from left to right, starting with the highest order term. an=10n-17 Multiply 10 by each term inside the parentheses. a110=1100-17 Subtract 17 from 1100 to get 1083. a110=1083